The t statistic for a test of

?0:?=11

??:?>11


based on n = 17 observations has the value t = 1.72.

(a) What are the degrees of freedom for this statistic?

(b) Using the appropriate table in your formula packet, bound the p-value as closely as possible:
< p-value <
Answer #1

Ho M = 11 HA. M711 n=17, t =1072 = 17 -1 = 1 Degree of foodom of (63) P-value 1- P(te 1:72) (Right tail) = 1 - 0.94764 test 0